7. The Inquiry Club at Jefferson University has compiled a book that exposes the private lives of many of the professors on campus. Economics majors in the club estimate that total revenue from sales of the book is given by the equation. TR = 120Q - 0.1Q3 Initially, the price is set at $71.60. To maximize total revenue, should the price be increased or decreased? 


1
Expert's answer
2022-02-28T16:07:56-0500

"TR = 120Q - 0,1Q^3"

"MR=TR'=(120Q - 0,1Q^3)'=120-0.3Q^2"

P=MR=MC

"120-0.3Q^2=71.60"

Q=12.70

"Prifit=TR-TC=71.60\times12.70-(120\times12.70-0.1(12.70)^3))=-409.842"


To maximize total revenue the price should be increased